OpenStudy (v-sync):

What the measure of each interior angle of a regular octagon? give answer and show work.

OpenStudy (v-sync):

I NEED THE FORMULA!!!!!!!!!!

OpenStudy (v-sync):

@CderaAlien @Google @Advert @Apple @Atsie @brainlesslonelygirl @cuanchi @Kevin @KittiWitti1

OpenStudy (ipurple):

12

OpenStudy (ipurple):

wait no

OpenStudy (ipurple):

theres 4 so 4x180=720

OpenStudy (v-sync):

There's 4 what?

OpenStudy (ipurple):

angels

OpenStudy (ipurple):

its a rectangular octagon

OpenStudy (v-sync):

What? but it's an octagon

OpenStudy (ipurple):

so i can be divided into 4 triangles

OpenStudy (ipurple):

that means the sum of all angels is 4*180=720

OpenStudy (ipurple):

degrees

OpenStudy (ipurple):

we dont use the mid angles

OpenStudy (v-sync):

But I don't need sum of all angles, I need measure of EACH angle.

OpenStudy (ipurple):

well that woud be 180 then

OpenStudy (ipurple):

because if 180 times 4 the number of angels is 720. so if you divided 720 bye 4 the number of angles you get 180

OpenStudy (v-sync):

That's not possible, 180 degree would be a straight line.

OpenStudy (ipurple):

4*180-360=720-360=360. and if we mesaure each interior angle in a square we get 90 degrees